OpenStudy (gorica):

I need help in finding inverse Laplace transformation of e^(-5s)/(2s^2+s+2)

hartnn (hartnn):

still need help ?

OpenStudy (gorica):

yes

hartnn (hartnn):

do you know the identity \(\huge L^{-1}[e^{-as}F(s)]=f(t-a)u(t-a)\) ?

OpenStudy (gorica):

yes

hartnn (hartnn):

so, you first need laplace inverse of 1/(2s^2+s+2) right ?

OpenStudy (gorica):

yes, I think I have it.

hartnn (hartnn):

ohh...so you were able to find laplace inverse of 1/(2s^2+s+2) ??

OpenStudy (gorica):

yes. It's this \[\frac{ 2 }{ \sqrt{15} }e ^{-t/4}\sin(\frac{ \sqrt{15} }{ 4 }t)\]

OpenStudy (gorica):

I hope :D

OpenStudy (sirm3d):

\[\checkmark\]

OpenStudy (gorica):

and this is where I don't know what to do next

hartnn (hartnn):

yes, thats correct, now just replace, t by 't-5' and add u(t-5) in the end. using that formula....

hartnn (hartnn):

\(\huge \frac{ 2 }{ \sqrt{15} }e ^{(-t+5)/4}\sin(\frac{ \sqrt{15} }{ 4 }(t-5))u(t-5)\)

OpenStudy (gorica):

thank you :)

hartnn (hartnn):

welcome ^_^